OpenStudy (mathmath333):

\(\large\tt \color{black}{a^3=abc+2,\\b^3=abc-3,\\c^3=abc+3,\\find~~the~~lowest~~number~~of~~a,~~b,~~c~~\\if~~a,~~b,~and~c~~are~~three~~real~~numbers}\)

OpenStudy (mathmath333):

\(\large\tt \color{black}{a. 3^{\frac{1}{3}}\\b. 9^{\frac{1}{3}}\\c. 2\\d.none~ of ~these}\)

OpenStudy (mathmath333):

the lowest number of a,b,c

OpenStudy (amistre64):

does solving for abc and equating help us any?

OpenStudy (mathmath333):

but i need some easy method solving through options i mean

OpenStudy (mathmath333):

is the question clear?

OpenStudy (mathmath333):

a,b and c are three real numbers and i have to find the lowest of them

OpenStudy (mathmath333):

solving the equations only spins my head

OpenStudy (kirbykirby):

Does it work if you solve everything for "abc", and then equate those equations together?

OpenStudy (mathmath333):

no i tried that but i cant found the right path,it only works when u substitute 3^(1/3) in b

OpenStudy (mathmath333):

have u find anything

OpenStudy (kirbykirby):

I've been going pretty much round in circles with these equations. I did find one equation that seemed useful, but how to solve it I wasn't sure. Wolfram popped out \(b=3^{1/3}\) and \(b=-\dfrac{3^{2/3}}{2^{1/3}}\).

OpenStudy (kirbykirby):

which was the solution to \(3b^3+11=3b(b^3+5)^{1/3}(b^3+6)^{1/3}+2\) but not sure how to solve that

OpenStudy (mathmath333):

ok' if i have to put option in variable a,b and c and to save time' how do i decide here in which of variable a,b or c should i substitute options for trial and error

OpenStudy (mathmath333):

@aum

OpenStudy (mathmath333):

@cj49

OpenStudy (aum):

If a, b and c are all positive, then b must be the smallest number because b^3 = abc - 3 (largest number being subtracted from abc). The next higher number would be a and the highest number would be c. Equate each answer choice to b and see if that works.

OpenStudy (mathmath333):

yes b works for option \(\large\tt \color{black}{a. 3^{\frac{1}{3}}}\)

OpenStudy (aum):

Does it work for the other two options?

OpenStudy (mathmath333):

no it doesnt work

OpenStudy (mathmath333):

thanks very much for reviewing this

OpenStudy (aum):

You are welcome.

OpenStudy (aum):

Wait a minute.... Is the 1/3 in the answer choices exponents?

OpenStudy (mathmath333):

yes of cos

OpenStudy (aum):

Then the lowest of \(\Large \sqrt[3]{3}, ~~\sqrt[3]{9}, ~~ 2\) is \(\Large \sqrt[3]{3}\) and that is the answer because with b set to \(\Large \sqrt[3]{3}\) you can solve for a and c and get real values which will be higher than b making b the lowest number.

OpenStudy (aum):

Initially the answer choices looked like 3 and 1/3; 9 and 1/3 rather than exponents.

OpenStudy (mathmath333):

yes it gives both values of \(\large\tt \color{black}{a~~and~~c}\)

OpenStudy (aum):

alright.

OpenStudy (mathmath333):

a3 = (abc + 2), b3 = (abc - 3), c3 = (abc + 3). Multiply all three equations and keep abc = x. x3 = (x + 2)(x - 3)(x + 3) --> 2x2 - 9x - 18 = 0 --> (2x + 3)(x - 6) = 0 --> x = -3/2 or x = 6. Taking abc = 6 and noting that b is the smallest number b3 = 6 - 3 = 3 --> b = 3^(1/3)